Two firms Intel and AMD in the CPU market have combined demand given by Q = 100 - 2 P. Their total costs are given by TC Intel = 2 Q Intel + Q^2 Intel and TC AMD = 2 Q AMD + Q^2AMD. If they cannot successfully collude and instead produce where the market price equals marginal cost, the market price will be:

Answers: 66

66.67

34

32

98
Answer #1

Date Page 8. QI-Q Intel O A = Q AMD 100-2P I + QA = 100-2P P= 100-(QA+Q=) 2 TCI 2 WIt Q I MC I = 2+2QI P = 100 (A+& I 2 100-